Nuance Company had net credit sales for the year of $500,000. Nuance estimates that 2 percent of its net credit sales will never be collected.

1.

Prepare the entry to record Nuance’s bad debt expense for the year.

2.

Nuance had accounts receivable of $100,000 at the end of the year. Show how the net accounts receivable balance would be reported on the balance sheet. Assume that the allowance for doubtful accounts had a beginning balance of zero.

3.

Why is A/R shown at net rather than just showing the full amount?


1
Expert's answer
2021-07-16T09:37:43-0400

Solution:

1.). Bad debt expense = Net credit sales x uncollectable 2%

= 500,000 "\times" 2"\%" % = 10,000

Bad debt expense = 10,000

Journal entry to record Nuance’s bad debt expense for the year:

Dr. Bad debt expense                                       10,000

             Cr. Allowance for doubtful accounts                      10,000

(To record bad debt expense that is unrecoverable)

 

2.). Net accounts receivable balance will be reported as follows in the balance sheet:

Allowance for doubtful debts = 10,000 + 0 = 10,000

 

Accounts receivable                            100,000

Less: Allowance for doubtful debts (10,000)

Net accounts receivable                        90,000

 

3.). This is because the company will not be able to collect all of its accounts receivable. A such a provision is made, which is deducted from the total accounts receivable figure. The estimated net realizable figure is the one that is shown on the financial statements to allow the users of the financial information to see the actual amount that the company believes will be collected from the total accounts receivable, hence the reason for showing A/R as a net figure.